OpenStudy (anonymous):

can 192 be evenly dived by 3

OpenStudy (anonymous):

help me please and thank you

OpenStudy (anonymous):

Yes it can. Because 192 ÷ 3 = 64

OpenStudy (anonymous):

192 divided by 3 = 64

OpenStudy (alexandervonhumboldt2):

1+9+2=12 12/3=4 so yes it can be divided you count need really to divide it